13 yellow balls and 14 red balls are placed in an urn. Two balls are then drawn in succession without replacement. What is the probability that the first ball drawn is a red ball if the second ball drawn is yellow?

a) 1.0000

b) 0.5556

c) 0.5385

d) 0.4815

e) 0.1378

f) None of the above.
Answer #1

P(2nd ball is yellow) = P(1st ball is red and 2nd ball is yellow) + P(1st ball is yellow and 2nd ball is yellow)

                                 = 14/27 * 13/26 + 13/27 * 12/26

                                 = 0.4815

P(1st ball is red | 2nd ball is yellow) = P(1st ball is red and 2nd ball is yellow) / P(2nd ball is yellow)

                                                          = (14/27 * 13/26) / 0.4815

                                                          = 0.5384

Option-C) 0.5385